OpenStudy (princesssleelee):

Describe how to transform the quantity of the fifth root of x to the seventh power, to the third powerinto an expression with a rational exponent. Make sure you respond with complete sentences.

OpenStudy (princesssleelee):

\[\left( \sqrt[5]{x^7} \right)^3\]

OpenStudy (princesssleelee):

@SolomonZelman

OpenStudy (prettygirl_shynice):

@princesssleelee u got it

OpenStudy (princesssleelee):

kind of, can you put it in relation to this specific problem?

OpenStudy (prettygirl_shynice):

what you mean

OpenStudy (princesssleelee):

you wrote for example "the root" would you be able to write the root in that spot? or is that how it is written for the answer to the problem?

OpenStudy (princesssleelee):

@Prettygirl_shynice

OpenStudy (prettygirl_shynice):

hold on brb

OpenStudy (princesssleelee):

ok

OpenStudy (prettygirl_shynice):

@princesssleelee u got it

OpenStudy (princesssleelee):

i think so, is that the complete answer?

OpenStudy (princesssleelee):

@Prettygirl_shynice

OpenStudy (princesssleelee):

@563blackghost

OpenStudy (princesssleelee):

@AlexandervonHumboldt2

OpenStudy (prettygirl_shynice):

im here

OpenStudy (princesssleelee):

i got x^21/5

OpenStudy (prettygirl_shynice):

u think it is alright

OpenStudy (alexandervonhumboldt2):

yeah correct

OpenStudy (princesssleelee):

what is @AlexandervonHumboldt2

OpenStudy (prettygirl_shynice):

u did good @princesssleelee

OpenStudy (princesssleelee):

i used a calculator i do not know how to do it step by step

OpenStudy (prettygirl_shynice):

oh

OpenStudy (princesssleelee):

That is what I am asking.

OpenStudy (alexandervonhumboldt2):

eute use power rule first: \[\huge \sqrt[a]{x^b}=x^{\frac{ a }{ b }}\]

OpenStudy (alexandervonhumboldt2):

woowps

OpenStudy (alexandervonhumboldt2):

i meant x^b/a not a/b

OpenStudy (princesssleelee):

Are you able to write it out in words please

OpenStudy (princesssleelee):

@mthompson440

OpenStudy (princesssleelee):

@tkhunny

OpenStudy (anonymous):

Hope this works for u.. At work so I had to rush without explanation

OpenStudy (princesssleelee):

i appreciate it, however if you could just write what you did, i would be completely satisfied and give you a medal :) lol

OpenStudy (princesssleelee):

@SolomonZelman

OpenStudy (princesssleelee):

@AlexandervonHumboldt2

OpenStudy (anonymous):

is it this\[\left( \sqrt[5]{x^7} \right)^3\]?

OpenStudy (princesssleelee):

the picture posted is perfect however idk what was done and word form is what im looking for

OpenStudy (princesssleelee):

yes that is it @satellite73

OpenStudy (anonymous):

oh i just saw the screen shot that looks good

OpenStudy (princesssleelee):

yes i just do not know the steps taken and i need them in a word form

OpenStudy (anonymous):

multiply seven times three get twenty one that goes in the numerator of the exponent the root, five, goes in the denominator of the exponent

OpenStudy (princesssleelee):

thats it?

OpenStudy (anonymous):

that is it, yes, i am sure you could have said that for yourself right/?

OpenStudy (princesssleelee):

i guess so lol.